OpenStudy (anonymous):

if a^n = (squareroot)a^3, then n = ?

OpenStudy (anonymous):

it means \[a ^{n}=\sqrt{a ^{3}}\]

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

yeahh. how did you get that?

OpenStudy (anonymous):

then a ^n = a^(3/2) so a = 3/2

OpenStudy (anonymous):

yeah 3/2=a

OpenStudy (anonymous):

just think as the SQRT sign at 1/2 when you raise that to a power SQRT(A)^n=A^(n)(1/2)=A^n/2

OpenStudy (anonymous):

ohh.oh. got it. *sheepish grin*

OpenStudy (anonymous):

i dont know why i cant do an equation its not letting me but nSQRT(a^m) whenre the n is on the front of the sqrt sign=a^m/n

OpenStudy (anonymous):

\[\sqrt[a]{n^b}=n^b/a\]